OpenStudy (kaylak12345):

Write an algebraic expression that models the word phrase. the sum of the product of a number x and 14​, and 8

OpenStudy (kittiwitti1):

the sum of, addition product, multiply

OpenStudy (kittiwitti1):

the sum of (blah, blah, product blah) and 8 means you add 8 all that product stuff means you multiply that.

OpenStudy (kaylak12345):

so is it (14+8)*x or 14x+8

OpenStudy (kaylak12345):

so 14x+8

OpenStudy (kittiwitti1):

\(\color{blue}{\text{Originally Posted by}}\) kittiwitti1 the sum of (blah, blah, product blah) \(\color{red}{\text{and 8 means you add 8}}\) all that product stuff means you multiply that. \(\color{blue}{\text{End of Quote}}\)

OpenStudy (kittiwitti1):

Yeah.

OpenStudy (kaylak12345):

ok good im doing corrections my grade needs to be pulled up

OpenStudy (kittiwitti1):

Alright good luck ♣

OpenStudy (kaylak12345):

can you help me on a few more

OpenStudy (kaylak12345):

2(m-6)+3<7

OpenStudy (kittiwitti1):

in words?

OpenStudy (kaylak12345):

solve the inequality

OpenStudy (kittiwitti1):

treat it like an equation with the equals sign EXCEPT when you divide by a negative number then you have to flip the sign around

OpenStudy (kaylak12345):

-8?

OpenStudy (kittiwitti1):

Why negative 8?

OpenStudy (kaylak12345):

one my choices are either 8 or negative and two positive 8 would make the answer 7 on that side too

OpenStudy (kittiwitti1):

Wait... you want the answer to m?

OpenStudy (kaylak12345):

that's how you would solve the inequality lol

OpenStudy (kittiwitti1):

\[2(m-6)+3\lt7\]\[2(m-6)\lt7-3\]\[m-6\lt\frac{4}{2}\]\[m\lt2+6\]\[m\lt8\]

OpenStudy (kaylak12345):

thats the answer

OpenStudy (kittiwitti1):

Yup

OpenStudy (kaylak12345):

3x-16+3x=4+6x-20 always sometimes or never true

OpenStudy (kittiwitti1):

what

OpenStudy (kittiwitti1):

I wouldn't know that one tbh

OpenStudy (kittiwitti1):

@mathmate

OpenStudy (kaylak12345):

hurry

OpenStudy (mathmate):

@kaylak12345 The question is badly worded. I would have put it as "Is the following statement always, sometimes, or never true? 3x-16+3x=4+6x-20 " Steps to answer this question: Simplify separately the left and right side of the equation. Three cases are possible: 1. left=right If the left-side comes out exactly equal to the right, then it is always true, because whatever you put the value of x, the two sides will remain equal. Example: 4x+3=4x+3, .... always true 2. left does not equal right, but the coefficients of x on each side are not equal, then it is sometimes true, because we can solve for x, and it is true ONLY when x equals this value. Example: 4x+2=3x+5 Solve for x: x=3. So the statement is true ONLY when x=3. 3. Coefficients of x are equal, but the constant terms are not. In this case, the statement can never be true. Example: 4x+3=4x+4 which when simplified, gives 3=4 which is never true. Simplify the given statement and find out which case it belongs to.

OpenStudy (mathmate):

@kaylak12345 Are you in a hurry? Are you doing a test? Please be informed that students are NOT allowed to ask for help at Open Study when doing tests, quizzes, or any other time limited evaluations.

OpenStudy (kaylak12345):

no im in a hurry because i have to go soon

OpenStudy (kaylak12345):

how do i solve this 4x-2=32x+19

OpenStudy (kaylak12345):

ill stay for 10 more minutes

OpenStudy (mathmate):

Is this the same question? If not, I suggest you post a new question.

OpenStudy (kaylak12345):

ok i will and i have another 30 minutes